Required information [The following information applies to the questions displayed below.] Cane Company manufactures two prod11. How many pounds of raw material are needed to make one unit of each of the two products? Beta Alpha Pounds of raw materia12. What contribution margin per pound of raw material is earned by each of the two products? (Round your answers to 2 decima13. Assume that Canes customers would buy a maximum of 85,000 units of Alpha and 65,000 units of Beta. Also assume that the14. Assume that Canes customers would buy a maximum of 85,000 units of Alpha and 65,000 units of Beta. Also assume that the15. Assume that Canes customers would buy a maximum of 85,000 units of Alpha and 65,000 units of Beta. Also assume that the
Answer #1

Answer 10.

Differential Analysis- Cane Company - Making Alpha (alt 1) or Buying Alph (Alt2)
Particulars Making Alpha (Alt 1) Buying Alpha (Alt 2) Financial advantage (Disadvantage) of buying (Alternative 2)
Costs:
Purchase Price (50000*$80) $0.00 $5,500,000 -$5,500,000
Direct material $1,650,000 $0.00 $1,650,000.00
Direct Labor $1,375,000 $0.00 $1,375,000.00
Variable manufacturing overhead $660000 $0.00 $660,000.00
Avoidable Fixed manufacturing Overhead $2247000 $0.00 $2247000
Total Cost $5,932,000.00 $5,500,000.00 $432000

Financial advantage =  $ 432,000

Answer 11.

Pounds of raw material needed to make one unit of:

Alpha=$30/5 = 6 pound per unit

Beta=$10/5 = 2 pound per unit

Answer 12.

Alpha beta
Sales revenue(per unit) $150 $105
Less variable cost per unit

$84

(30+25+12+17)

$53

(10+20+10+13)

Contribution margin $66 $52
Pounds of direct material needed to produce 1 unit 6 2
Contribution margin per pound $11 $26

Answer 13.

Optimal no. Of units to produce would be computed as:

Product pound per unit unit produced total pounds
Beta 2 65000 130000
Alpha 6 6000 36000

Answer 14.

The total contribution margin would be computed as follow:

Alpha Beta
No. Of units produced 6000 65000
Contribution margin per unit $66 $52
Total contribution margin $396,000 $3,380,000

Company total contribution margin would be $3,776,000 [396000+3380000]

Answer 15.

The maximum price per pound is computed as follows:

Alpha

Regular direct material cost per pound = $5

Contribution margin per pound of direct materials = 11

Maximum price to be paid per pound = 5 + 11 = $16

Because the company has satisfied all demand for Betas, it would use additional raw materials to produce Alphas.